A person can drive from town A to town B at a certain rate of speed in 6 hours. If he increases his speed by 10 kph, he can make the trip in 5 hours. How far is it from town A to town B?
Select the correct answer.
a. 300 km
b. 330 km
c. 320 km
d. 310 km

Answer by mananth(16949)

let speed be x kph
so distance = 6x
increased speed = (x+10) kph
time = 5 hours
distance = 5(x+10)
6x=5(x+10)
6x=50+5x
x=50
distance = 50*6
=300 km